### v1

# Past Movements-Guided Motion Representation Learning for Human Motion Prediction

Abstract
Human motion prediction based on 3D skeleton is a significant challenge in computer vision, primarily focusing on the effective representation of motion. In this paper, we propose a self-supervised learning framework designed to enhance motion representation. This framework consists of two stages: first, the network is pretrained through the self-reconstruction of past sequences, and the guided reconstruction of future sequences based on past movements. We design a velocity-based mask strategy to focus on the joints with large-scale moving. Subsequently, the pretrained network undergoes finetuning for specific tasks. Self-reconstruction, guided by patterns of past motion, substantially improves the model’s ability to represent the spatiotemporal relationships among joints but also captures the latent relationships between past and future sequences. This capability is crucial for motion prediction tasks that solely depend on historical motion data. By employing this straightforward yet effective training paradigm, our method outperforms existing state-of-the-art methods, reducing the average prediction errors by 8.8% across Human3.6M, 3DPW, and AMASS datasets. The code is available at https://github.com/JunyuShi02/PMG-MRL.
1 Introduction
3D skeleton-based motion prediction stands as a fundamental task in computer vision, with the objectives of generating consecutive and unseen future motion sequences based on past observations. This technique finds widespread application in domains such as human-computer interaction, autonomous driving, and motion analysis. The key to achieve accurate predictions of future motions lies in effectively modeling motion sequences, particularly the spatiotemporal representation of moving joints. However, due to the inherent properties of human kinematics, joints exhibit complex hierarchical rotational relationships. The motion trajectories of human joints also display a certain degree of randomness. Consequently, effectively capturing the spatiotemporal relationships between joints has perennially posed a challenging problem.
In the early stages, researchers utilized probabilistic models, such as Gaussian processes[38] and nonlinear Markov[19] models, to predict future motion changes based on historical motion sequences. However, these methods are computationally intensive and make strong assumptions about data distribution, which leads to models that are inadequate for real-world prediction scenarios. With the emergence of large-scale datasets[17, 37, 27] and advancements in hardware computing performance, motion prediction tasks started employing deep learning methods based on Recurrent Neural Networks (RNNs)[11, 16, 9, 14, 31]. RNN-based models effectively capture the dependencies between frames in a sequence, but the step-by-step generation approach results in the accumulation of prediction errors. The rise of Graph Neural Networks (GNNs) has brought feed-forward prediction paradigms into the motion prediction[26, 3, 10, 30, 34]. GNNs are well-suited for handling the non-Euclidean distribution properties of joints in spatial dimensions, enabling effective modeling of motion’s spatial information. Moreover, methods based on Transformers and Multi Layer Perceptrons (MLPs) have also achieved success in motion prediction tasks[1, 12, 2].
Overall, the mainstream paradigms for 3D skeleton-based motion prediction can be categorized into two main categories: 1) extracting features from historical motion sequences using an encoder and then generating predicted motions by a decoder[20, 11, 23] (as shown in Fig. 1(a)); 2) directly generating predicted motions using a feed-forward network[28, 26, 22] (as shown in Fig. 1(b)). However, the potential of self-supervised learning paradigms based on pretraining has not been sufficiently explored and applied in the field of motion prediction. This limits the ability of models to fully learn the spatiotemporal representations of motion sequences. AuxFormer[41] improves prediction performance by incorporating denoising tasks and masked feature prediction as auxiliary tasks during training (as shown in Fig. 1(c)). Nevertheless, parallel multi-task prediction encounters challenges such as interference with the main task (future motion prediction) and difficulties in balancing loss weights. Moreover, conducting masked feature prediction directly on the entire sequence lacks an exploration of the relationship between historical and future sequences.
In this paper, we propose a novel motion representation learning method for motion prediction tasks, which comprises two stages: pretraining and finetuning. In the pretraining stage, we apply masking to the joints of future motion sequences and utilize the complete past sequence features to guide the reconstruction of future sequences. This approach not only enhances the model’s reconstruction performance through past sequence features but also allows the model to learn the underlying relationship between historical and future sequences during the pretraining stage. Simultaneously, we also conduct self-reconstruction training based on masking on the past sequences. When applying the mask, we mask the joints with the largest motion changes between adjacent frames (i.e., the set of joints with the highest velocity). These joints serve as the central representation of the primary motion changes. Compared to random masking, this approach avoids learning motion patterns of joints with lower information content. For instance, in the ”Eating” action, motion changes are concentrated in the upper limb part, implying that we should focus on the representation learning of the upper limb part (including hand, head, and other joints), rather than the lower limb part with smaller changes.
Motion data contains both temporal and spatial dimensions. Given the outstanding performance of Transformers in temporal modeling tasks and spatial modeling tasks, as well as prior work on using Transformers for motion prediction tasks [41, 32, 1], we employ spatiotemporal Transformers to model motion sequences. Specifically, our method includes two components based on Transformers: the Past Motion Encoder (PME) and the Future Motion Predictor (FMP). The former comprises SpatioTemporal Attention (ST-Attention) modules, while the latter consists of ST-Attention modules and Past Motion Guided Attention (PMG-Attention) modules that fuse historical and future information. By pretraining the parameters of the PME and FMP, and finetuning them for the prediction task, our method significantly outperforms other state-of-the-art methods in both long-term and short-term predictions using a straightforward model architecture. It is worth noting that we observe a correlation between the performance improvement of our method and the increase in dataset size, which aligns with the characteristic of self-reconstruction learning requiring large-scale data.
The key contributions of our work can be summarized as follows:
Figure 1: (a) Prediction based on Encoder-Decoder paradigm[20] [23]. (b) Prediction based on Feed-forward paradigm[24] [28]. (c) Enhancing prediction based on auxiliary task[6] [41]. (d) Our Pretrain-Finetune prediction paradigm.
  • We propose a human motion representation learning method based on self-reconstruction pretraining paradigm for motion prediction tasks. This method helps the model learn the intrinsic spatiotemporal structure and features of motion sequences.
  • We use past motion features to guide the self-reconstruction of future sequences, which not only improves self-reconstruction performance but also helps the model understand the underlying relationship between past and future sequences in the motion prediction tasks.
  • We conduct extensive experiments and demonstrate that our method reduces the average prediction errors by 8.8% compared to state-of-the-art methods on the 3DPW, Human3.6M, and AMASS datasets.
2 Related Work
2.1 3D Skeleton-based Human Motion Prediction
RNN-based Methods
Early motion prediction methods employed probability model-based approaches[38, 19]. These methods performed poorly when dealing with complex motions. As RNNs can naturally process sequential data, researchers widely used RNNs and their variants for motion prediction tasks[11, 9, 14, 31]. EDR[9] maps motion sequences to a latent space using a nonlinear encoder and models temporal features using LSTM (Long Short-Term Memory Network). To avoid ambiguity when predicting long-term motions, VAE-LSTM[14] introduces the encoding of control signals into a variational inference framework trained to learn human motion manifolds. United[31] achieves smooth prediction by modeling first-order motion derivatives. To incorporate spatial information modeling, Structure-RNN[18] unfolds the spatiotemporal graph in the temporal dimension and decomposes it into a set of contributing factor components fed into RNN modeling. SkelNet[13], at each time frame, obtains local representations of human structure through special component layer branches.
Feed-forward Methods
The step-by-step modeling and generation approach of RNNs often leads to the issue of error accumulation. In recent times, many methods have adopted a feed-forward paradigm, utilizing GNN and Transformer models, to generate prediction sequences in a non-autoregressive manner[26, 3, 10, 30, 34]. These approaches leverage graph structures and attention computations that adapt to the non-Euclidean distribution properties of joints in space and time. LTD[28], HisRep[29], SGSN[22], and SPGSN[24] transform motion sequences to the frequency domain, modeling the trajectories of joints. DMGNN[20], MST-GNN[23], MSR-GCN[7], Sybio-GNN[21], and MMotion-Att[30] learn motion patterns at multiple levels by modeling human representations at different scales. DGMGNN[20], MANet[34], DD-GCN[40], and DSTD-GCN[10] model the dynamic characteristics of joints, adapting to the dynamic properties of motion. Transformer-based models utilize self-attention computations, which enhance the model’s capability to capture long-range dependencies between joints[1, 32, 41]. Among them, AuxFormer[41] achieves better results by incorporating denoising and reconstruction tasks to assist the prediction task. In addition, lightweight models based on MLP[2, 12] have also demonstrated success.
2.2 Self-supervised Pretraining
Self-supervised pretraining has brought about a revolution in the understanding and generation of language and images. This technique serves as one of the fundamental training methods of Large Language Models (LLMs), which is widely used in BERT[8], RoBERTa[25] and other models. For image processing tasks, recovering masked pixels/tokens is a common self-supervised pretraining method[15, 5]. MAE[15] proves that even when a substantial portion of image information is masked, the network can reconstruct a complete image based on a limited number of visible features. The self-learning paradigm of autoencoder in MAE is also widely used in multimodal[4], video processing[36], point cloud data processing[42] and other tasks.
3 Method
Figure 2: Detail of the model architecture. Our proposed model comprises Past Motion Encoder (PME) and Future Motion Predictor (FMP), where the former consists of $N\times$ SpatioTemporal Attention (ST-Attention), and the latter consists of $M$ ST-Attention and Past Motion Guided Attention (PMG-Attention). Here, the finetune stage is taken as an example to show the model structure (the pretrain stage only distinguishes the number of forward propagation of PME and incomplete input sequences).
3.1 Problem Background
The purpose of 3D skeleton-based human motion prediction is to generate the unseen and consecutive future motion based on past motion patterns. Assume a historical motion sequence consisting of $T$ frames in the Cartesian coordinate system, denoted as $\mathbf{X}^{P}=[\textbf{x}^{1},\textbf{x}^{2},...,\textbf{x}^{T}]\in\mathbb{R}^{T\times J\times K}$, where $J$ represents the number of joints and $K$ is the feature dimension (the initial value of $K$ is 3, representing the $x,y,z$ coordinates). Assume $\mathbf{X}^{F}=[\textbf{x}^{1},\textbf{x}^{2},...,\textbf{x}^{L}]\in\mathbb{R}^{L\times J\times K}$ is the future motion pose of $L$ frames. Our goal is to design and train a network $\mathcal{P}(\cdot)$ that takes $\mathbf{X}^{P}$ as input and $\mathbf{X}^{F}$ as output, i.e., $\mathbf{X}^{F}=\mathcal{P}(\mathbf{X}^{P},\pi_{\theta})$, where $\pi_{\theta}$ is the trainable parameters.
3.2 Model Architecture
As shown in Fig. 2, our proposed model consists of a Past Motion Encoder (PME) and a Future Motion Predictor (FMP). The PME aims to extract spatiotemporal features from the past motion sequence. The FMP aims to generate the future motion sequence based on the initial state (which consists of zero-valued inputs in motion prediction tasks) and past motion features.
Joint Embedder and Predict Head
Given an input sequence $\mathbf{X}$, we first use a single linear layer to embed each joint of each frame. To distinguish the spatiotemporal order information of the joints, we add trainable temporal and spatial position encoding to the motion embedding. This process can be represented as:
$\mathbf{H}=\mathbf{W}\mathbf{X}+\mathbf{b}+\mathbf{P}_{t}+\mathbf{P}_{s}$ (1)
where $\mathbf{W}\in\mathbb{R}^{K\times C},\mathbf{b}\in\mathbb{R}^{C}$ are the weights and biases of the linear layer, and $\mathbf{P}_{t},\mathbf{P}_{s}$ are the temporal and spatial position encoding, which are randomly initialized and trainable. $C$ denotes the number of feature channels. For the output of PME and FMP, a linear layer as the Predict Head is applied to obtain the 3D motion sequence.
Past Motion Encoder (PME)
We design PME to model the features of the past motion sequence. Human motion sequences contain two features: the change of joint position over time and the relationship between joints in space. We use the SpatioTemporal Attention (ST-Attention) to simultaneously model the spatiotemporal information of the joints. Specifically, PME contains $N$ cascaded ST-Attention modules, where each module first models the sequence dependencies in the temporal dimension and then models the joint relationships in the spatial dimension. Assuming the input to the $l^{th}$ layer of ST-Transformer is $\mathbf{H}^{l}$, the computation process can be expressed as:
$\mathbf{H}^{l+1}=\mathcal{F}_{attn}^{s}(\mathcal{F}_{attn}^{t}(\mathbf{H}^{l}))$ (2)
where $\mathcal{F}_{attn}^{s}(\cdot)$ / $\mathcal{F}_{attn}^{t}(\cdot)$ represents the spatial/temporal multi-head self-attention operation with the bottle neck.
Future Motion Predictor (FMP)
The purpose of FMP is to generate future motion sequences based on the initial state and the output from PME (i.e., past motion features). FMP comprises two modules: the ST-Attention module for modeling the sequence dependencies within itself and the Past Movement Guided Attention module (PMG-Attention). PMG-Attention applies cross-attention to incorporate past motion features into the current state, which can be mathematically represented as:
$\mathbf{H}^{l+1}=\mathcal{F}_{c-attn}^{s}(\mathcal{F}_{c-attn}^{t}(\mathbf{H}^{l},\mathbf{H}^{P}),\mathbf{H}^{P})$ (3)
where $\mathbf{H}^{P}$ is the past motion features output by PME. $\mathbf{H}^{l}$ is the input of $l^{th}$ PMG-Attention. $\mathcal{F}_{c-attn}^{s}(\cdot)$ / $\mathcal{F}_{c-attn}^{t}(\cdot)$ represents the spatial/temporal multi-head cross-attention calculation with the bottle neck. We provide more information about the ST-Attention and PMG-Attention in the supplementary materials (Sec. 7).
3.3 Past Movement Guided Pretraining
We apply the self-reconstruction pretraining paradigm to help the model learn the spatiotemporal representation of joints in motion sequences (as shown in the left part of Fig. 1(d)), which compromise three steps: 1) past motion reconstruction; 2) past motion encoding; 3) future motion reconstruction. We will introduce each step in the following.
Past Motion Reconstruction
Given the past motion sequence $\mathbf{X}^{P}\in\mathcal{R}^{T\times J\times K}$, we first calculate the motion velocity of each joint $\mathbf{V}=\mathbf{X}_{2:T}^{P}-\mathbf{X}_{1:T-1}^{P}\in\mathcal{R}^{(T-1)\times J\times K}$. We find the threshold $s$ in $\mathbf{V}$ such that the $r$of the values in $\mathbf{V}$ are greater than the threshold $s$. The mask of the past motion sequence $\mathbf{M}^{P}$ can be mathematically expressed as:
$\mathbf{M}^{P,(i,j)}=\left\{\begin{array}[]{ll}1&,\mathbf{V}^{(i,j)}\geq s\\ 0&,\mathbf{V}^{(i,j)}<s\end{array}\right.$ (4)
The masked past motion sequence can be computed by $\mathbf{X}_{masked}^{P}=\mathbf{X}^{P}\times\mathbf{M}^{P}$. Then, the joint embedders are applied to obtain the masked motion embeds $\mathbf{X}_{emb}^{P}$. Noting that we incorporate the trainable temporal and spatial position encoding into the motion embeds to learn the temporal order and spatial relationships of joints. Finally, we calculate the reconstructed motion sequence $\mathbf{X}_{rec}^{P}=\mathcal{F}_{PH}(\mathcal{F}_{PME}(\mathbf{X}_{emb}^{P}))$, where $\mathcal{F}_{PH}(\cdot)$ and $\mathcal{F}_{PME}(\cdot)$ are Predict Head and Position Motion Encoder.
Past Motion Encoding
This step aims to compute the features of complete past motion, which will be used to guide the future motion reconstruction and integrate the hidden relationship between the past sequence and the future sequence into the network. Given the complete input sequence $\mathbf{X}^{P}$, we first apply joint embedder to obtain the motion embeds $\mathbf{X}_{emb}^{P}$. Then, the past motion features can be computed by $\mathbf{H}^{P}=\mathcal{F}_{PME}(\mathbf{X}_{emb}^{P})$.
Future Motion Reconstruction
Given the future motion sequence $\mathbf{X}^{F}\in\mathcal{R}^{L\times J\times K}$, we first use the same method for masking and embedding the future motion sequence as for processing the past motion sequence. The masked future motion embeds, denoted as $\mathbf{X}_{emb}^{F}$ and the complete past motion features $\mathbf{H}^{P}$ are the input of FMP. In the PMG-Attention of FMP, past motion features guide the reconstruction of masked future motion through cross-attention mechanism. Finally, the reconstructed motion sequence can be expressed as: $\mathbf{X}_{rec}^{F}=\mathcal{F}_{PH}(\mathcal{F}_{FMP}(\mathbf{X}_{emb}^{F},\mathbf{H}^{P})))$.
3.4 Fine-tuning for Motion Prediction
In the finetune phase, we use the complete past motion sequence as input to the PME to get past motion features (as shown in the right part of Fig. 1(d)). In FMP, we use the initial state initialized with the value of 0 (but retain the position encoding) as input, and output the complete future sequence guided by the past motion features. Given the past motion sequence $\mathbf{X}^{P}$, this process can be represented as:
$\mathbf{H}^{P}$ $=\mathcal{F}_{JE}(\mathcal{F}_{PME}(\mathbf{X}^{P}))$ (5)
$\mathbf{H}^{F}$ $=\mathcal{F}_{JE}(\mathcal{F}_{FMP}(\mathbf{H}_{0}^{F},\mathbf{H}^{P}))$ (6)
$\mathbf{X}^{F}$ $=\mathcal{F}_{PH}(\mathbf{H}^{F})$ (7)
where $\mathbf{H}_{0}^{F}$ is the initial state initialized with the value of 0. $\mathcal{F}_{JE}$ represents the joint embedder.
3.5 Loss Function
Loss Function of Pretrain Stage
Pretrain stage performs past sequence reconstruction and future sequence reconstruction. Given the past and future reconstructed results $\hat{\mathbf{X}}_{rec}^{P}$ and $\hat{\mathbf{X}}_{rec}^{F}$, and the corresponding ground-truth (i.e. complete sequence) $\mathbf{X}_{rec}^{P}$ and $\mathbf{X}_{rec}^{F}$, the loss function can be represented as follow:
$\begin{split}\mathcal{L}_{pretrain}&=\frac{1}{TJ}\sum_{t=1}^{T}\sum_{j=1}^{J}||\mathbf{X}_{rec}^{P,(t,j)}-\hat{\mathbf{X}}_{rec}^{P,(t,j)}||_{F}^{2}\\ &+\frac{\alpha}{LJ}\sum_{l=1}^{L}\sum_{j=1}^{J}||\mathbf{X}_{rec}^{F,(l,j)}-\hat{\mathbf{X}}_{rec}^{F,(l,j)}||_{F}^{2}\end{split}$ (8)
where $\alpha$ is the hyperparameter to balance the loss weight.
Loss Function of finetune Stage
For finetune stage, given the predicted sequence $\hat{\mathbf{X}}$ and the ground-truth $\mathbf{X}$, the loss function can be defined as:
$\mathcal{L}_{finetune}=\frac{1}{LJ}\sum_{l=1}^{L}\sum_{j=1}^{J}||\mathbf{X}_{j}^{l}-\hat{\mathbf{X}}_{j}^{l}||_{F}^{2}$ (9)
Table 1: Comparisons of MPJPEs between our proposed method with other state-of-the-art methods for short-term prediction on Human3.6. We show the results of detailed actions such as walking and eating. The best results are highlighted in bold, and the second best results is underlined.
Motion Walking Eating Smoking Discussion
Milliseconds 80 160 320 400 80 160 320 400 80 160 320 400 80 160 320 400
DMGNN[20] 17.3 30.7 54.6 65.2 11.0 21.4 36.2 43.9 9.0 17.6 32.1 40.3 17.3 34.8 61.0 69.8
MSR[7] 12.2 22.7 38.6 45.2 8.4 17.1 33.0 40.4 8.0 16.3 31.3 38.2 12.0 26.8 57.1 69.7
Mixer[2] 12.0 23.0 42.4 51.3 7.8 16.5 34.3 43.0 7.4 15.2 30.3 37.6 10.7 25.0 56.9 71.1
PGBIG[26] 10.2 19.8 34.5 40.3 7.0 15.1 30.6 38.1 6.6 14.1 28.2 34.7 10.0 23.8 53.6 66.7
SPGSN[24] 10.1 19.4 34.8 41.5 7.1 14.8 30.5 37.9 6.7 13.8 28.0 34.6 10.4 23.8 53.6 67.1
siMLPe[12] 11.9 19.3 31.9 36.9 10.2 16.5 30.7 37.3 10.4 16.4 29.7 35.9 12.7 25.3 55.1 68.1
DSTD-GC[10] 11.1 22.4 38.8 45.2 7.0 15.5 31.7 39.2 6.6 14.8 29.8 36.7 10.0 24.4 54.5 67.4
AuxFormer[41] 8.9 16.9 30.1 36.1 6.4 14.0 28.8 35.9 5.7 11.4 22.1 27.9 8.6 18.8 38.8 49.2
CIST-GCN[33] 11.8 23.4 40.5 46.5 6.7 14.8 29.8 36.8 7.3 15.6 31.0 38.0 10.2 23.7 52.3 65.3
Ours 8.6 16.5 31.2 37.7 6.1 13.7 29.3 36.8 5.3 10.3 19.7 25.2 7.7 16.3 31.7 40.9
Motion Direction Greeting Phoning Posing
Milliseconds 80 160 320 400 80 160 320 400 80 160 320 400 80 160 320 400
DMGNN[20] 13.1 24.6 64.7 81.9 23.3 50.3 107.3 132.1 12.5 25.8 48.1 58.3 15.3 29.3 71.5 96.7
MSR[7] 8.6 19.7 43.3 53.8 16.5 37.0 77.3 93.4 10.1 20.7 41.5 51.3 12.8 29.4 67.0 85.0
Mixer[2] 7.5 18.4 43.9 55.3 16.4 37.1 80.2 98.2 9.1 19.5 42.4 53.6 11.5 27.0 64.9 83.5
PGBIG[26] 7.2 17.6 40.9 51.5 15.2 34.1 71.6 87.1 8.3 18.3 38.7 48.4 10.7 25.7 60.0 76.6
SPGSN[24] 7.4 17.2 40.0 50.3 14.6 32.6 70.6 86.4 8.7 18.3 38.7 48.5 10.7 25.3 59.9 76.5
siMLPe[12] 11.0 19.4 41.6 52.1 16.2 33.7 70.9 86.3 11.9 20.3 39.7 48.9 13.9 27.3 62.0 78.6
DSTD-GC[10] 6.9 17.4 40.1 51.7 14.3 33.5 72.2 87.3 8.5 19.2 40.3 49.9 10.1 25.4 60.6 77.3
AuxFormer[41] 6.8 17.0 40.3 51.6 13.5 31.3 69.2 85.4 7.9 17.3 37.4 47.2 8.8 19.1 39.2 51.0
CIST-GCN[33] 7.3 18.1 43.6 55.3 13.7 31.0 65.7 79.9 8.6 18.5 39.3 49.6 9.6 23.7 57.7 75.0
Ours 6.3 15.8 39.8 51.4 10.8 26.4 62.7 79.8 7.5 17.1 37.9 48.0 8.2 16.9 32.8 43.5
Table 2: Comparisons of MPJPEs between our proposed method with other state-of-the-art methods for short- and long- term prediction on Human3.6M. We show the average result of all actions. The best results are highlighted in bold.
Motion Average Param(M)
Milliseconds 80 160 320 400 560 1000 -
DMGNN[20] 17.0 33.6 65.9 79.7 103.0 137.2 4.82
MSR[7] 12.1 25.6 51.6 62.9 81.1 114.2 6.30
Mixer[2] 11.1 24.0 51.5 64.0 83.5 116.5 0.03
PGBIG[26] 10.3 22.7 47.4 58.5 76.9 110.3 1.80
SPGSN[24] 10.4 22.3 47.1 58.3 77.1 108.8 5.66
siMLPe[12] 13.4 23.7 47.4 58.0 77.9 111.7 0.14
DSTD-GC[10] 10.4 23.3 48.8 59.8 77.8 111.0 0.18
AuxFormer[41] 9.5 20.6 43.4 54.1 75.3 107.0 1.00
CIST-GCN[33] 10.5 23.2 47.9 59.0 77.2 110.3 0.35
GCNext [39] 9.3 21.5 45.5 56.4 74.7 108.7 -
Ours 8.8 19.3 41.1 51.8 70.0 101.5 1.66
4 Experiments
4.1 Dataset
Human3.6M
Human3.6M[17] is a large-scale, accurately annotated 3D human pose dataset that captures the motions of 11 professional actors in daily life scenarios using motion capture technology, totaling over 3.6 million pose instances. Following previous works[26, 24], only 22 joints are used, focusing on those that exhibit significant movement, and all sequences are down-sampled to 25 fps.
The 3D Poses in the Wild (3DPW)
3DPW[37] stands out as a pioneering dataset in motion related tasks, primarily due to its focus on capturing human motion in natural settings outside of the typical laboratory environment. This dataset provides a rich collection of video sequences captured at a frame rate of 30 frames per second, which includes diverse activities such as ’Walking’, ’Running’, and interacting with various objects.
The Archive of Motion Capture as Surface Shapes (AMASS)
AMASS[27] consolidates 15 existing motion capture datasets into a single, standardized framework. By converting these diverse datasets into a common representation, AMASS provides over 40 hours of meticulously annotated motion data. Following the previous works[34, 40], we discarded the detailed joints of the hand and sampled the sequence to 24fps.
Figure 3: Visualization results of short-term prediction (a) and reconstruction (b). The blue lines represent predicted/reconstructed motions, and the red lines represent ground-truths. We map the results to 2D space for display.
4.2 Experimental Settings
For the PME and the FMP, the number of modules (denoted as $N,M$) are both set to 3. The number of feature channels remains consistent at 128 throughout the network architecture. In the attention calculation, we employed 8 heads, with each head comprising 32 feature channels. During the pretraining phase, the loss weight (denoted as $\alpha$) is set to 1, indicating an equal contribution of loss from both past and future sequence reconstruction. We provide more implementation details in the supplementary materials (Sec. 8).
4.3 Comparison to State-of-the-art Methods
We compare our method with the recent state-of-the-art methods, including DMGNN (CVPR’20)[20, 30, 1, 7, 2, 26, 24, 12, 10, 41, 40, 33, 39]. To ensure fair comparisons, we re-evaluated some methods on the entire test set to align the testing standard using their official pretrained models, or retrain the model (Only for the cases where the official pre-training model is not provided) using their default hyperparameters on the Human3.6M.
Comparisons on Human3.6M
We use MPJPE (Mean Per Joint Position Error) as the evaluation metric, which is widely used in motion prediction tasks[26, 24, 41]. Tab. 1 shows the quantitative comparisons of short-term motion prediction on Human3.6M dataset, which aims to predict the poses within 400 milliseconds. Detailed actions are listed. As can be seen, our method outperforms the others on most actions, especially the ”Smoking”, ”Greeting”, and ”Posing”. Tab. 2 displays the average results of short - and long-term predictions for all actions on Human3.6M. Our approach effectively reduces prediction errors by 5.4%/10.2%/9.7%/8.2%/6.3%/6.6% at 80ms/160ms/320ms/400ms/560ms/1000ms.
Table 3: Comparisons of MPJPEs between our proposed method with other state-of-the-art methods for both short-term and long-term prediction on AMASS. The best results are highlighted in bold.
Milliseconds 80 160 320 400 560 720 880 1000
LTD-10-10[28] 10.3 19.3 36.6 44.6 61.5 75.9 86.2 91.2
LTD-10-25[28] 11.0 20.7 37.8 45.3 57.2 65.7 71.3 75.2
HisRep[29] 11.3 20.7 35.7 42.0 51.7 58.6 63.4 67.2
Mao W, et al.[30] 11.0 20.3 35.0 41.2 50.7 57.4 61.9 65.8
siMLPe[12] 10.8 19.6 34.3 40.5 50.5 57.3 62.4 65.7
DD-GCN[40] 10.4 19.1 33.6 39.8 49.3 56.5 61.3 64.6
Ours 8.6 15.3 27.7 33.2 42.0 48.3 53.0 56.0
Comparisons on AMASS
Tab. 3 shows the comparisons on AMASS dataset. Our method achieves a reduction in prediction errors of 17.3%/19.9%/17.6%/16.6%/14.8%/14.5%/13.5%/13.3% at 80ms/160ms/320ms/400ms/560ms/720ms/880ms/1000ms. We found that our method exhibits the most significant improvement on this dataset. This can be attributed to the larger data size of AMASS compared to Human3.6M and 3DPW. The self-reconstruction pretraining enables our model to learn a more comprehensive representation of motion on such a large dataset.
Table 4: Comparisons of MPJPEs between our proposed method with other state-of-the-art methods for short- and long-term prediction on 3DPW. The best results are highlighted in bold.
Average MPJPE
Milliseconds 100 200 400 600 800 1000
DMGNN[20] 17.80 37.11 70.38 94.12 109.67 123.93
HisRep[29] 15.88 35.14 66.82 93.55 107.63 114.75
MSR-GCN[7] 15.70 33.48 65.02 93.81 108.15 116.31
STSGCN[35] 18.32 37.79 67.51 92.75 106.65 112.22
PGBIG[26] 17.66 35.32 67.83 89.60 102.59 109.41
SPGSN[24] 15.39 32.91 64.54 91.62 103.98 111.05
AuxFormer[41] 14.21 30.04 58.50 89.45 100.78 107.45
Ours 13.85 29.11 57.06 87.35 97.93 103.37
Comparisons on 3DPW
The 3DPW dataset comprises data collected from outside the laboratory setting, which are widely used to verify the power of the model in real scenarios. As shown in Tab. 4, our approach performed best across all frames, showcasing its robustness in handling complex interactive scenarios.
Table 5: Comparison of Motion Classification Using Network-Output Features
w/o Network AuxFormer[41] Ours
Accuracy 25.6% 80.8% 95.0%
Representation Capability
To validate the representation learning capability of our proposed method, we first extract motion features from the outputs of the trained model (outputs from the penultimate layer). Subsequently, we train a simple linear classifier to classify these features. As shown in Tab. 5, our method significantly improves classification accuracy compared to methods that do not involve feature extraction via neural networks and AuxFormer[41].
4.4 Visualization of Prediction and Reconstruction
Fig. 3 (a) presents the visualization of prediction results. We conducted comparisons with PGBIG[26] and Mixer[2] on challenging samples. In the ’Posing’ and ’Discussion’ actions, our model successfully captures significant upward/downward movements of the arms. This can be attributed to the utilization of past motion guidance, which enables the model to comprehend the influence of previous motion patterns on future sequences. We also provide reconstruction results with a mask rate of 75% in Fig. 3 (b). Similar to the property of mask-based pretraining[15], even with a high mask rate of 75% for joints, the model can still accurately reconstruct the exact location of them.
5 Ablation Study
In this section, we explore the impact of components of our network, including the mask strategy, pretraining paradigm. We provide studies on model architecture in the supplementary materials. All experiments were carried out on the Human3.6M dataset.
Mask Strategy
Tab. 6 presents the results of two cases: A, without utilizing the velocity-based mask strategy (i.e. randomly mask 75% of the joints), and F, which represents the complete model. It can be seen that using the velocity-based mask reduced the average prediction errors by 2.3%. Tab. 6 also shows the comparisons of different mask rates (ranging from 0% to 75%), represented by B, C, D, F. The high mask rate facilitates the model to learn the space-time representation during the self-reconstruction process.
Effects of Pretraining Paradigm
We show the results of pretraining the model by denoising in Tab. 6 E. Compared with the self-reconstruction, denoising pretraining method lacks the exploration of the spatiotemporal relationship between joints.
Table 6: Impact of mask strategy and pretraining paradigm. $r$ is the mask rate. ’denoise’ means we conduct self-supervised pretraining by denoising.
ID 80ms 160ms 320ms 400ms
A random mask 9.0 19.7 42.4 52.8
B $r=0\%$ 9.6 21.9 46.2 56.8
C $r=25\%$ 9.3 20.2 42.9 53.7
D $r=50\%$ 9.1 20.0 42.3 52.6
E denoise 9.2 20.4 42.9 53.8
F complete ($r=75\%$) 8.8 19.3 41.1 51.8
6 Conclusion
In this paper, we propose a novel motion prediction paradigm, which consists two stages: 1) Past motion guided future motion self-reconstruction for learning the spatio-temporal representation of human joints (i.e. pretrain stage). 2) Train the network for the motion prediction task (i.e. finetune stage). Despite the simplicity of the model structure, our approach achieves the best results on three representative datasets and is highly extensible.
Supplementary Material
Table 7: Comparisons of MPJPEs between our proposed method with other state-of-the-art methods for short-term prediction on Human3.6. We show the results across all actions such as walking and eating. The best results are highlighted in bold, and the second best results is underlined
Motion Walking Eating Smoking Discussion Direction
Milliseconds 80 160 320 400 80 160 320 400 80 160 320 400 80 160 320 400 80 160 320 400
DMGNN[20] 17.3 30.7 54.6 65.2 11.0 21.4 36.2 43.9 9.0 17.6 32.1 40.3 17.3 34.8 61.0 69.8 13.1 24.6 64.7 81.9
MSR[7] 12.2 22.7 38.6 45.2 8.4 17.1 33.0 40.4 8.0 16.3 31.3 38.2 12.0 26.8 57.1 69.7 8.6 19.7 43.3 53.8
Mixer[2] 12.0 23.0 42.4 51.3 7.8 16.5 34.3 43.0 7.4 15.2 30.3 37.6 10.7 25.0 56.9 71.1 7.5 18.4 43.9 55.3
PGBIG[26] 10.2 19.8 34.5 40.3 7.0 15.1 30.6 38.1 6.6 14.1 28.2 34.7 10.0 23.8 53.6 66.7 7.2 17.6 40.9 51.5
SPGSN[24] 10.1 19.4 34.8 41.5 7.1 14.8 30.5 37.9 6.7 13.8 28.0 34.6 10.4 23.8 53.6 67.1 7.4 17.2 40.0 50.3
siMLPe[12] 11.9 19.3 31.9 36.9 10.2 16.5 30.7 37.3 10.4 16.4 29.7 35.9 12.7 25.3 55.1 68.1 11.0 19.4 41.6 52.1
DSTD-GC[10] 11.1 22.4 38.8 45.2 7.0 15.5 31.7 39.2 6.6 14.8 29.8 36.7 10.0 24.4 54.5 67.4 6.9 17.4 40.1 51.7
AuxFormer[41] 8.9 16.9 30.1 36.1 6.4 14.0 28.8 35.9 5.7 11.4 22.1 27.9 8.6 18.8 38.8 49.2 6.8 17.0 40.3 51.6
CIST-GCN[33] 11.8 23.4 40.5 46.5 6.7 14.8 29.8 36.8 7.3 15.6 31.0 38.0 10.2 23.7 52.3 65.3 7.3 18.1 43.6 55.3
Ours 8.6 16.5 31.2 37.7 6.1 13.7 29.3 36.8 5.3 10.3 19.7 25.2 7.7 16.3 31.7 40.9 6.3 15.8 39.8 51.4
Motion Greeting Phoning Posing Purchases Sitting
Milliseconds 80 160 320 400 80 160 320 400 80 160 320 400 80 160 320 400 80 160 320 400
DMGNN[20] 23.3 50.3 107.3 132.1 12.5 25.8 48.1 58.3 15.3 29.3 71.5 96.7 21.4 38.7 75.7 92.7 11.9 25.1 44.6 50.2
MSR[7] 16.5 37.0 77.3 93.4 10.1 20.7 41.5 51.3 12.8 29.4 67.0 85.0 14.8 32.4 66.1 79.6 10.5 22.0 46.3 57.8
Mixer[2] 16.4 37.1 80.2 98.2 9.1 19.5 42.4 53.6 11.5 27.0 64.9 83.5 13.4 30.7 65.6 80.0 9.8 20.9 45.6 57.3
PGBIG[26] 15.2 34.1 71.6 87.1 8.3 18.3 38.7 48.4 10.7 25.7 60.0 76.6 12.5 28.7 60.1 73.3 8.8 19.2 42.4 53.8
SPGSN[24] 14.6 32.6 70.6 86.4 8.7 18.3 38.7 48.5 10.7 25.3 59.9 76.5 12.7 28.6 61.0 74.4 9.3 19.4 42.2 53.6
siMLPe[12] 16.2 33.7 70.9 86.3 11.9 20.3 39.7 48.9 13.9 27.3 62.0 78.6 15.5 29.3 59.3 72.0 13.2 21.9 44.5 55.7
DSTD-GC[10] 14.3 33.5 72.2 87.3 8.5 19.2 40.3 49.9 10.1 25.4 60.6 77.3 12.7 29.7 62.3 75.8 8.8 19.3 42.9 54.3
AuxFormer[41] 13.5 31.3 69.2 85.4 7.9 17.3 37.4 47.2 8.8 19.1 39.2 51.0 11.9 28.0 61.8 76.3 8.7 19.0 42.1 53.3
CIST-GCN[33] 13.7 31.0 65.7 79.9 8.6 18.5 39.3 49.6 9.6 23.7 57.7 75.0 13.3 30.2 63.0 77.3 8.9 19.4 42.3 53.6
Ours 10.8 26.4 62.7 79.8 7.5 17.1 37.9 48.0 8.2 16.9 32.8 43.5 11.4 27.6 62.6 77.6 8.5 19.2 44.3 56.7
Motion Sitting Down Taking Photo Waiting Walking Dog Walking Together
Milliseconds 80 160 320 400 80 160 320 400 80 160 320 400 80 160 320 400 80 160 320 400
DMGNN[20] 15.0 32.9 77.1 90.3 13.6 29.0 46.0 58.8 12.2 24.2 59.6 77.5 47.1 93.3 160.1 171.2 14.3 26.7 50.1 63.2
MSR[7] 16.1 31.6 62.5 76.8 9.9 21.0 44.6 56.3 10.7 23.1 48.3 59.2 20.7 42.9 80.4 93.3 10.6 20.9 37.4 43.9
Mixer[2] 15.3 30.5 62.7 78.0 9.0 19.9 44.7 56.9 9.9 22.0 49.7 62.5 21.1 43.9 84.7 101.2 9.9 20.4 39.3 47.3
PGBIG[26] 13.9 27.9 57.4 71.5 8.4 18.9 42.0 53.3 8.9 20.1 43.6 54.3 18.8 39.3 73.7 86.4 8.7 18.6 34.4 41.0
SPGSN[24] 14.2 27.7 56.7 70.7 8.8 18.9 41.5 52.7 9.2 19.8 43.1 54.1 - - - - 8.9 18.2 33.8 40.9
siMLPe[12] 18.0 29.8 57.7 71.2 13.1 21.3 42.9 53.6 11.9 21.5 44.4 55.0 21.8 40.6 74.6 87.0 11.9 19.8 34.2 39.8
DSTD-GC[10] 14.1 28.0 57.3 71.2 8.4 18.8 42.0 53.5 8.7 20.2 44.3 55.3 19.6 41.8 77.6 90.2 9.1 19.8 36.3 42.7
AuxFormer[41] 13.5 27.6 57.7 72.2 8.2 18.4 41.5 53.0 8.2 18.5 41.2 52.2 17.1 36.5 70.4 83.0 7.8 15.9 30.2 37.0
CIST-GCN[33] 14.1 29.8 57.3 69.8 8.2 18.4 40.6 51.8 8.6 19.4 43.5 54.8 20.0 41.4 73.7 85.1 9.6 20.3 38.2 45.6
Ours 10.1 17.2 30.2 39.2 7.6 17.8 42.0 54.1 7.8 18.0 41.6 53.2 19.1 41.1 80.6 95.2 7.3 15.3 30.8 37.9
In this supplementary material, we provide more information that could not be included in the main manuscript because of space limit, including explanation of ST-Attention and PMG-Attention (in Sec. 7), implementation details (in Sec. 8), complete results for testing on Human3.6M dataset (in Sec. 9), ablation study on model architecture (in Sec. 10), and visualization of feature map (in Sec. 11).
7 Detailed Explanation of ST-Attention and PMG-Attention
In our model architecture for human motion prediction, both the Past Motion Encoder (PME) and the Future Motion Predictor (FMP) modules rely on attention mechanisms: SpatioTemporal Attention (ST-Attention) and Past Movement Guided Attention (PMG-Attention). These mechanisms are pivotal for capturing and integrating spatiotemporal dependencies in motion sequences.
ST-Attention
ST-Attention combines temporal and spatial attention mechanisms to model joint relationships over time effectively. This attention mechanism is utilized within PME to encode past motion features.
The temporal attention mechanism operates on the sequence dependencies over time in the joint positions:
$\mathbf{Q}^{t},\mathbf{K}^{t},\mathbf{V}^{t}=\mathcal{F}^{t}_{QKV}\left(\mathbf{H}^{l}\right)\$ (10)
$\mathbf{Z}^{t}=\operatorname{softmax}\left(\frac{\mathbf{Q}^{t}\left(\mathbf{K}^{t}\right)^{\mathrm{T}}}{\sqrt{d_{k}}}\right)\mathbf{V}^{t}$ (11)
where $\mathbf{H}^{l}$ represents the input at the $l^{th}$ layer, $\mathbf{Q}^{t},\mathbf{K}^{t}$, and $\mathbf{V}^{t}$ are queries, keys, and values, respectively, derived from $\mathbf{H}^{l}$, and $d_{k}$ is the dimension of keys.
The spatial attention mechanism focuses on modeling joint relationships in space:
$\mathbf{Q}^{s},\mathbf{K}^{s},\mathbf{V}^{s}=\mathcal{F}^{s}_{QKV}\left(\mathbf{Z}^{t}\right)$ (12)
$\mathbf{Z}^{s}=\operatorname{softmax}\left(\frac{\mathbf{Q}^{s}\left(\mathbf{K}^{s}\right)^{\mathrm{T}}}{\sqrt{d_{k}}}\right)\mathbf{V}^{s}$ (13)
where $\mathbf{Q}^{s},\mathbf{K}^{s}$, and $\mathbf{V}^{s}$ are queries, keys, and values for spatial attention, respectively, similarly derived from $\mathbf{Z}^{t}$. $\mathbf{Z}^{s}$ is the output of ST-Attention.
PMG-Attention
PMG-Attention is employed within the Future Motion Predictor (FMP) module to incorporate past motion features into the prediction of future motion sequences. Unlike ST-Attention which focuses on the past motion sequence itself, PMG-Attention applies cross-attention between the current state and the past motion features:
$\mathbf{Q}^{t},\mathbf{K}^{t},\mathbf{V}^{t}=\mathcal{F}^{t}_{QKV}\left(\mathbf{H}^{l}\right),\mathcal{F}^{t}_{QKV}\left(\mathbf{H}^{P}\right)$ (14)
$\mathbf{Z}^{t}=\operatorname{softmax}\left(\frac{\mathbf{Q}^{t}\left(\mathbf{K}^{t}\right)^{\mathrm{T}}}{\sqrt{d_{k}}}\right)\mathbf{V}^{t}$ (15)
$\mathbf{Q}^{s},\mathbf{K}^{s},\mathbf{V}^{s}=\mathcal{F}^{s}_{QKV}\left(\mathbf{Z}^{t}\right),\mathcal{F}^{s}_{QKV}\left(\mathbf{H}^{P}\right)$ (16)
$\mathbf{Z}^{s}=\operatorname{softmax}\left(\frac{\mathbf{Q}^{s}\left(\mathbf{K}^{s}\right)^{\mathrm{T}}}{\sqrt{d_{k}}}\right)\mathbf{V}^{s}$ (17)
where $\mathbf{Q}^{t}$ and $\mathbf{Q}^{s}$ are obtained from the hidden state of future motions, and the $\mathbf{K}^{t},\mathbf{V}^{t},\mathbf{K}^{s},\mathbf{V}^{s}$ are obtained from the features of past motion.
8 Implementation Details
We deployed our network using an NVIDIA RTX4090 GPU (with PyTorch 2.0). For training, we set the batch size to 24 and utilized the Adam optimizer with an initial learning rate of 0.0005. To enhance convergence, we implemented a cosine annealing learning rate schedule. These training parameters were consistently applied across all datasets in our experiments. We ensured that the GPU memory usage remained below 6GB, enabling our network to be seamlessly deployed on various hardware platforms with ease.
9 Complete Results on Human3.6M
We provide the full results of comparisons on the Human3.6M dataset in Tab. 7 as a supplement to Tab. 1. Despite the simplicity of our network structure, we achieved the best results on most actions.
10 Ablation Study on Model Architecture
Table 8: Ablation study of model architecture
ID 80ms 160ms 320ms 400ms
A add 9.4 20.1 42.7 53.5
B connect 9.2 20.1 42.7 53.7
C parallel attn. 9.1 20.1 42.3 52.9
D complete 8.8 19.3 41.1 51.8
Way of Fusing Past Motion and Future Motion
In the default setting, we use cross-attention to integrate features of past and future motions. We experimented with other fusion methods, with results shown in Tab. 8: A) Direct addition of features from past and future actions; B) Concatenation of both in the feature dimension, followed by a Linear layer to reduce the number of channels back to the original value. We observed that cross-attention is the best choice because it explicitly computes the temporal and spatial interactions between the past and future motion.
Table 9: Impact of model size. $d$ is the feature dimension. $l$ is the number of layers.
ID 80ms 160ms 320ms 400ms Param
E $d=64$ 9.3 20.2 43.3 54.4 0.67M
F $d=128$ 9.1 19.8 42.0 52.9 1.00M
G $l=2$ 9.2 20.2 42.6 53.3 1.16M
H $l=4$ 8.8 19.3 41.6 52.3 2.15M
D complete 8.8 19.3 41.1 51.8 1.66M
Figure 4: Visualization of attention feature maps.
Spatiotemporal Attention
In the default setting, we follow the AuxFormer[41] of computing temporal and spatial attention sequentially. Results of parallel computation of temporal and spatial attention are presented in Tab. 8 C, showing worse performance.
Model Size
In Tab. 9, we show the impact of different model sizes (including feature dimension $d$ in attention layer and number of layers $l$) on parameter count and results. Although reducing the model size can decrease the number of parameters, it can also lead to decreased performance. We observe that the default setting ($d$=256, $l$=3) represents the optimal balance between parameter count and performance.
11 Visualization of Attention Feature Map
For the example of the Discussion action, we present the averaged attention feature maps over 512 samples. As shown in Fig. 4, sub figure (a) and (b) display the temporal self-attention and spatial self-attention feature maps from different layers in PME, sub figure (c) and (d) show the temporal cross-attention and spatial cross-attention feature maps from different layers in FMP.
In temporal cross-attention, the model’s first and third layers tend to allocate attention to the later frames, as these frames have a greater influence on future sequences. In the second layer, the model distributes attention to both later and earlier frames, highlighting the potential impact of historical sequences on future actions. In spatial cross-attention, the model identifies significant correlations between the movement patterns of the ’thorax’, ’head’, ’left elbow’, and ’right elbow’ joints and the prediction of future actions.